# HopRAG: Multi-Hop Reasoning for Logic-Aware Retrieval-Augmented Generation

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) systems often struggle with imperfect retrieval, as traditional retrievers focus on lexical or semantic similarity rather than logical relevance. To address this, we propose HopRAG, a novel RAG framework that augments retrieval with logical reasoning through graph-structured knowledge exploration. During indexing, HopRAG constructs a passage graph, with text chunks as vertices and logical connections established via LLM-generated pseudo-queries as edges. During retrieval, it employs a retrieve-reason-prune mechanism: starting with lexically or semantically similar passages, the system explores multi-hop neighbors guided by pseudo-queries and LLM reasoning to identify truly relevant ones. Extensive experiments demonstrate HopRAG’s superiority, achieving 76.78% higher answer accuracy and 65.07% improved retrieval F1 score compared to conventional methods. The repository is available at https://github.com/LIU-Hao-2002/HopRAG.
@@
Hao Liu1†, Zhengren Wang1,2†, Xi Chen3, Zhiyu Li2∗ Feiyu Xiong2, Qinhan Yu1, Wentao Zhang1∗ 1Peking University 2Center for LLM, Institute for Advanced Algorithms Research, Shanghai 3Huazhong University of Science and Technology {liuhao_2002,wzr,yuqinhan}@stu.pku.edu.cn       xichenai@hust.edu.cn {lizy, xiongfy}@iaar.ac.cn       wentao.zhang@pku.edu.cn
@@
[1.5em]1.5em1em $\dagger$ The first two authors have equal contributions.
* Corresponding Author.
@@ 1 Introduction
From a logical structure perspective, existing RAG systems can be mainly categorized into three types: Non-structured RAG simply adopts sparse or dense retrievers. The retrieval is only based on keyword matching or semantic vector similarity, but fails to capture the logical relations between user queries and passages. Tree-structured RAG (Sarthi et al., 2024; Chen et al., 2023; Fatehkia et al., 2024) focuses on the hierarchical logic of passages within a single document, but ignores relations beyond the hierarchical structure or across documents. Additionally, it introduces redundant information across different levels. Graph-structured RAG (Soman et al., 2024; Kang et al., 2023; Edge et al., 2024a; Guo et al., 2024) models logical relations in the most ideal form by constructing knowledge graphs (KGs) to represent documents, where entities are vertices and their relations are edges. However, the reliance on predefined schemas limits the flexible expressive capability (Li et al., 2024); constructing and updating knowledge graphs is challenging and prone to errors or omissions (Edge et al., 2024a); the triplet format of knowledge necessitates extra textualization or fine-tuning to improve LLMs’ understanding (He et al., 2024).
@@
Based on these observations, we propose HopRAG, an innovative graph-structured RAG system. At indexing phase, we construct a graph-structured knowledge index with passages as vertices and logic relations as directed edges. Specifically, the passages are connected by pseudo-queries generated by query simulation and edge merging operations. For example, as demonstrated in Figure 2, the pseudo-query "What does the frog do?" connects the raiser passage and the solver passage, as the pivot for logical hops. During retrieval, we employ reasoning-augmented graph traversal, following a three-step paradigm of retrieval, reasoning, and pruning. This process searches for truly relevant passages within the multi-hop neighborhood of indirectly relevant passages, guided by both the index structure and LLM reasoning.
@@
  • Extensive experiments confirm the effectiveness of HopRAG. The retrieve-reason-prune mechanism achieves over 76.78% higher answer metric and 65.07% higher retrieval F1 score compared to conventional RAG approaches. Several ablation studies provide more valuable insights.
@@ 3.2 Graph-Structured Index
We construct a graph-structured index $G=(\mathcal{V},\mathcal{E})$ where we store each passage in a vertex to get vertex set $\mathcal{V}$ and capture the logical relations between passages to establish directed edge set =$\mathcal{E}=\{\langle v_{i},e_{i,j},v_{j}\rangle|v_{i},v_{j}\in\mathcal{V}\}$ for multi-hop reasoning ($\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$). To identify the relations between passages and establish the graph-structured index, we utilize query simulation to dig out the logical relations between passages and leverage textual similarity for efficient edge merging.
@@
To fully explore the logical relations between passages and bridge the inherent gap between queries and passages (Wang et al., 2024c), we generate a series of pseudo-queries for each passage to explicitly capture the logical relations. Specifically, we adopt LLM to generate two groups of pseudo-queries for each passage: (1) $m$ out-coming questions $Q_{i}^{+}=\bigcup_{1\leq j\leq m}\{q_{i,j}^{+}\}$ that originate from the passage but cannot be answered by itself; (2) $n$ in-coming questions $Q_{i}^{-}=\bigcup_{1\leq j\leq n}\{q_{i,j}^{-}\}$ whose answers are within the passage. The prompts are in Appendix A.5.
We extract keywords from $Q_{i}^{+}$ and $Q_{i}^{-}$ using named entity recognition $\text{NER}(\cdot)$ for sparse representation, and embed these questions into semantic vectors using an embedding model $\text{EMB}(\cdot)$ for dense representation. This results in sparse representations $K_{i}^{+}=\bigcup_{1\leq j\leq m}\{k_{i,j}^{+}\}$ and $K_{i}^{-}=\bigcup_{1\leq j\leq n}\{k_{i,j}^{-}\}$, and dense representations $V_{i}^{+}=\bigcup_{1\leq j\leq m}\{v_{i,j}^{+}\}$ and $V_{i}^{-}=\bigcup_{1\leq j\leq n}\{v_{i,j}^{-}\}$. We further define out-coming triplets as $r_{i,j}^{+}:=(q_{i,j}^{+},k_{i,j}^{+},v_{i,j}^{+})$ and in-coming triplets $r_{i,j}^{-}:=(q_{i,j}^{-},k_{i,j}^{-},v_{i,j}^{-})$. Each passage $p_{i}$ is a vertex $v_{i}$, featured with out-coming triplets $R_{i}^{+}=\bigcup_{1\leq j\leq m}\{r_{i,j}^{+}\}$ and in-coming triplets $R_{i}^{-}=\bigcup_{1\leq j\leq n}\{r_{i,j}^{-}\}$.
@@
Given the out-coming and in-coming triplets, we match paired triplets via hybrid retrieval and establish directed edges between passages. For each out-coming triplet $r_{s,i}^{+}$ of source vertex $v_{s}$, the most matching in-coming triplet $r_{t^{*},j^{*}}^{-}$ is determined as follows:
$\text{SIM}(r_{s,i}^{+},r_{t,j}^{-})$ $=\frac{\frac{|k_{s,i}^{+}\cap k_{t,j}^{-}|}{|k_{s,i}^{+}\cup k_{t,j}^{-}|}+\frac{v_{s,i}^{+}\cdot v_{t,j}^{-}}{||v_{s,i}^{+}||\cdot||v_{t,j}^{-}||}}{2}$ (1)
@@
We then build the directed edge $\langle v_{s},e_{s,t^{*}},v_{t^{*}}\rangle$ with aggregated attribution, where $e_{s,t^{*}}:=({q_{t^{*},j^{*}}^{-}},{k_{t^{*},j^{*}}^{-}\cup k_{s,i}^{+}},{v_{t^{*},j^{*}}^{-}})$.
@@ 3.3 Reasoning-Augmented Graph Traversal
For more accurate and complete responses, HopRAG’s retrieval strategy leverages the reasoning ability of LLMs to explore the neighborhood of probably indirectly relevant passages based on the logical relations in the graph structure so as to hop to relevant ones. As shown in Algorithm 1, by reasoning over the questions on out edges $e_{i,j}$ of a current vertex $v_{i}$ and then choosing to hop to the most promising vertex $v_{j}$, we realize reasoning-augmented graph traversal for better retrieval performance.
@@
To start the local search over the graph for query $q$, we first use $\text{NER}(\cdot)$ and $\text{EMB}(\cdot)$ to get the keywords $k_{q}$ and vector $v_{q}$ of $q$, which will be used for hybrid retrieval to match $top_{k}$ similar $\langle v_{i},e_{i,j},v_{j}\rangle$ edges, following Equation 1. With each head vertex $v_{j}$ from these edges we initialize a context queue $C_{queue}$ for breadth-first local search (Voudouris et al., 2010).
@@
To fully exploit the logical relations over the graph and hop from indirectly relevant vertices to relevant ones, we introduce breadth-first local search which utilizes the LLM to choose the most appropriate neighbor for each $v_{j}$ in $C_{queue}$ to append to the tail of the queue. Specifically, for each $v_{j}$ in $C_{queue}$ in each round of hop, we leverage LLM to reason over all the questions from its out edges to choose one $e_{j,k}$ with the question which the LLM regards as the most helpful for answering $q$ and append the tail vertex $v_{k}$ to $C_{queue}$.
After hopping from each vertex in the current $C_{queue}$ we can expand the context with $top_{k}$ logically more relevant vertices, from which we continue to conduct the next round of hop. By conducting $n_{hop}$ rounds of hop, we realize reasoning-augmented graph traversal that expands the context to $(n_{hop}+1)\times top_{k}$ vertices. We use a counter $C^{\prime}$ to track the number of arrivals of each vertex during the traversal.
@@
To avoid including too many intermediate vertices during the traversal, we introduce a novel metric Helpfulness $H(\cdot)$ that integrates similarity and logic to re-rank and then prune the traversal context $C^{\prime}$ with arrival counts. We calculate $H_{i}$ following Equation 2 for each $v_{i}$ in $C^{\prime}$ and keep the $top_{k}$ vertices with the highest $H_{i}$, where hybrid textual similarity $\text{SIM}(v_{i},q)$ follows Equation 1 to calculate the average lexical and semantic similarity between the passage in $v_{i}$ and query $q$; and $\text{IMP}(v_{i},C^{\prime})$ is defined as the normalized number of arrivals of $v_{i}$ in $C^{\prime}$ during traversal (Equation 3). With $H_{i}$ for each $v_{i}$ in $C^{\prime}$, we prune $C^{\prime}$ by keeping $top_{k}$ vertices with the highest $H$ and get final context $C$.
$H_{i}=\frac{\text{SIM}(v_{i},q)+\text{IMP}(v_{i},C^{\prime})}{2}$ (2)
$\text{IMP}(v_{i},C^{\prime})=\frac{C^{\prime}[v_{i}]}{\sum_{v_{j}\in C^{\prime}}C^{\prime}[v_{j}]}$ (3)
Input: $q$, $top_{k}$, $n_{hop}$
@@ 1 $v_{q}$leftarrow EMB($q$);
3 $C_{queue}\leftarrow$ Retrieve($v_{q}$, $k_{q}$);
4 $C^{\prime}\leftarrow$Counter$(C_{queue})$;
@@ 4.1 Experimental Setups
We collect several multiple-hop QA datasets to evaluate the performance of HopRAG. We use HotpotQA dataset (Yang et al., 2018), 2WikiMultiHopQA dataset (Ho et al., 2020) and MuSiQue dataset (Trivedi et al., 2022). Following the same procedure as (Zhang et al., 2024), we obtain 1000 questions from each validation set of these three datasets. See Appendix A.2 for details.
@@
We compare HopRAG with a variety of baselines: (1) unstructured RAG - sparse retriever BM25 (Robertson and Zaragoza, 2009a) (2) unstructured RAG - dense retriever BGE (Xiao et al., 2024; Karpukhin et al., 2020) (3) tree-structured RAG - RAPTOR (Sarthi et al., 2024) (4) tree-structured RAG - SiReRAG (Zhang et al., 2024) (5) graph-structured RAG - GraphRAG (Edge et al., 2024a) with the local search function (6) graph-structured RAG - HippoRAG (Gutiérrez et al., 2025). For structured RAG baselines, we follow the same setting as previous work (Zhang et al., 2024).
@@
We adopt exact match (EM) and F1 scores on answers to measure the QA performance of different methods, which focus on the accuracy between a generated answer and the corresponding ground truth. We do not use retrieval metrics to compare all the methods because some baseline methods like SiReRAG (Zhang et al., 2024) and RAPTOR (Sarthi et al., 2024) create new candidates (e.g., summary nodes) in the retrieval pool and it would be unfair to use retrieval metrics to compare the methods. But we report both the answer metrics and the retrieval metrics in the ablation study when comparing the effects of different hyperparameters and reasoning models on HopRAG. See Appendix A.3 for more metric details.
@@
We adopt the BGE embedding model for semantic vectors at 768 dimensions. To avoid the loss of semantic information caused by chunking at a fixed size, we adopt the same chunking methods utilized in the original datasets respectively. GPT-4o-mini serves as both the model generating in-coming and out-coming questions when constructing the graph index, and the reasoning model for graph traversal. We use two reader models GPT-4o and GPT-3.5-turbo to generate the response given the context with the fixed size of 20 retrieval candidates and at most $n_{hop}=4$. See Appendix A.4 for more setting details.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) systems often struggle with imperfect retrieval, as traditional retrievers focus on lexical or semantic similarity rather than logical relevance. To address this, we propose HopRAG, a novel RAG framework that augments retrieval with logical reasoning through graph-structured knowledge exploration. During indexing, HopRAG constructs a passage graph, with text chunks as vertices and logical connections established via LLM-generated pseudo-queries as edges. During retrieval, it employs a retrieve-reason-prune mechanism: starting with lexically or semantically similar passages, the system explores multi-hop neighbors guided by pseudo-queries and LLM reasoning to identify truly relevant ones. Experiments on multiple multi-hop benchmarks demonstrate that HopRAG’s retrieve-reason-prune mechanism can expand the retrieval scope based on logical connections and improve final answer quality.
@@
Hao Liu1†, Zhengren Wang1,2†, Xi Chen3, Zhiyu Li2∗, Feiyu Xiong2, Qinhan Yu1, Wentao Zhang1∗ 1Peking University 2Center for LLM, Institute for Advanced Algorithms Research, Shanghai 3Huazhong University of Science and Technology {liuhao_2002,wzr,yuqinhan}@stu.pku.edu.cn       xichenai@hust.edu.cn {lizy, xiongfy}@iaar.ac.cn       wentao.zhang@pku.edu.cn
@@
[1.5em]1.5em1em $\dagger$ Equal contribution; * Corresponding author.
@@ 1 Introduction
From a logical structure perspective, existing RAG systems can be mainly categorized into three types: Non-structured RAG simply adopts sparse or dense retrievers. The retrieval is only based on keyword matching or semantic vector similarity, but fails to capture the logical relations between user queries and passages. Tree-structured RAG (Sarthi et al., 2024; Chen et al., 2023; Fatehkia et al., 2024) focuses on the hierarchical logic of passages within a single document, but ignores relations beyond the hierarchical structure or across documents. Further, it introduces redundant information across different levels. Graph-structured RAG (Soman et al., 2024; Kang et al., 2023; Edge et al., 2024a; Guo et al., 2024) models logical relations in the most ideal form by constructing knowledge graphs (KGs) to represent documents, where entities are vertices and their relations are edges. However, the reliance on predefined schemas limits the flexible expressive capability (Li et al., 2024); constructing and updating knowledge graphs are challenging and prone to errors or omissions (Edge et al., 2024a); the triplet format of knowledge necessitates extra textualization or fine-tuning to improve LLMs’ understanding (He et al., 2024).
@@
Based on these observations, we propose HopRAG, an innovative graph-structured RAG system. At indexing phase, we construct a graph-structured knowledge index with passages as vertices and logic relations as directed edges. Specifically, the passages are connected by pseudo-queries generated by query simulation and edge merging operations. For example, as demonstrated in Figure 2, the pseudo-query "Why does the princess kiss the frog?" connects the raiser passage and the solver passage, as the pivot for logical hops. During retrieval, we employ reasoning-augmented graph traversal, following a three-step paradigm of retrieval, reasoning, and pruning. This process searches for truly relevant passages within the multi-hop neighborhood of indirectly relevant passages, guided by both the index structure and LLM reasoning.
@@
  • Extensive experiments confirm the effectiveness of HopRAG. The retrieve-reason-prune mechanism achieves over 36.25% higher answer metric and 20.97% higher retrieval F1 score compared to conventional information retrieval approaches. Several ablation studies provide more valuable insights.
@@ 3.2 Graph-Structured Index
We construct a graph-structured index $G=(\mathcal{V},\mathcal{E})$ where the vertex set $\mathcal{V}$ consists of vertices storing all the passages and the directed edge set =$\mathcal{E}=\{\langle v_{i},e_{i,j},v_{j}\rangle|v_{i},v_{j}\in\mathcal{V}\}\subset\mathcal{V}\times\mathcal{V}$ is established based on the logical relations between passages for multi-hop reasoning. To establish $G$, we utilize Query Simulation to identify the logical relations and leverage textual similarity for efficient Edge Merging.
@@
To identify the logical relations between passages, we generate a series of pseudo-queries for each passage, and use them to explore the passage’s relations with the others and bridge the inherent gap between user-queries and passages (Wang et al., 2024c). Specifically, we adopt LLM to generate two groups of pseudo-queries for each passage $p_{i}$: (1) $m$ out-coming questions $Q_{i}^{+}=\bigcup_{1\leq j\leq m}\{q_{i,j}^{+}\}$ that originate from this passage but cannot be answered by itself; (2) $n$ in-coming questions $Q_{i}^{-}=\bigcup_{1\leq j\leq n}\{q_{i,j}^{-}\}$ whose answers are within the passage. As demonstrated in Figure 2, for the toy passage "Rose is the princess in the story The Frog Prince", one in-coming question might be "What is the name of the princess?" and one out-coming question might be "How is the frog connected to Rose?". The prompts are in Appendix A.5.
We extract keywords from $Q_{i}^{+}$ and $Q_{i}^{-}$ using named entity recognition $\text{NER}(\cdot)$ for sparse representation, and embed these questions into semantic vectors using an embedding model $\text{EMB}(\cdot)$ for dense representation. This results in sparse representations $K_{i}^{+}=\bigcup_{1\leq j\leq m}\{k_{i,j}^{+}\}$ and $K_{i}^{-}=\bigcup_{1\leq j\leq n}\{k_{i,j}^{-}\}$, and dense representations $V_{i}^{+}=\bigcup_{1\leq j\leq m}\{\mathbf{v}_{i,j}^{+}\}$ and $V_{i}^{-}=\bigcup_{1\leq j\leq n}\{\mathbf{v}_{i,j}^{-}\}$. We further define out-coming triplets as $r_{i,j}^{+}:=(q_{i,j}^{+},k_{i,j}^{+},\mathbf{v}_{i,j}^{+})$ and in-coming triplets $r_{i,j}^{-}:=(q_{i,j}^{-},k_{i,j}^{-},\mathbf{v}_{i,j}^{-})$. Each passage $p_{i}$ is stored inside a vertex $v_{i}$, featured with its out-coming triplets $R_{i}^{+}=\bigcup_{1\leq j\leq m}\{r_{i,j}^{+}\}$ and in-coming triplets $R_{i}^{-}=\bigcup_{1\leq j\leq n}\{r_{i,j}^{-}\}$.
@@
Given the out-coming and in-coming triplets, we match paired triplets via hybrid retrieval and establish directed edges between the corresponding passages. For each out-coming triplet $r_{s,i}^{+}$ of source vertex $v_{s}$, the most matching in-coming triplet $r_{t^{*},j^{*}}^{-}$ is determined as follows:
$\text{SIM}(r_{s,i}^{+},r_{t,j}^{-})$ $=\frac{\frac{|k_{s,i}^{+}\cap k_{t,j}^{-}|}{|k_{s,i}^{+}\cup k_{t,j}^{-}|}+\frac{\mathbf{v}_{s,i}^{+}\cdot\mathbf{v}_{t,j}^{-}}{||\mathbf{v}_{s,i}^{+}||\cdot||\mathbf{v}_{t,j}^{-}||}}{2}$ (1)
@@
We then build the directed edge $\langle v_{s},e_{s,t^{*}},v_{t^{*}}\rangle$ with aggregated features, where $e_{s,t^{*}}:=({q_{t^{*},j^{*}}^{-}},{k_{t^{*},j^{*}}^{-}\cup k_{s,i}^{+}},{\mathbf{v}_{t^{*},j^{*}}^{-}})$.
@@ 3.3 Reasoning-Augmented Graph Traversal
For more accurate and complete responses, HopRAG’s retrieval strategy leverages the reasoning ability of LLM to explore the neighborhood of probably indirectly relevant passages and hop to relevant ones based on the logical relations in the graph structure. As shown in Algorithm 1, by reasoning over the questions on out edges $e_{i,j}$ of a current vertex $v_{i}$ and then choosing to hop to the most promising vertex $v_{j}$, we realize reasoning-augmented graph traversal for better retrieval performance.
@@
To start the local search over the graph for query $q$, we first use $\text{NER}(\cdot)$ and $\text{EMB}(\cdot)$ to get the keywords $k_{q}$ and vector $\mathbf{v}_{q}$ of $q$, which will be used for hybrid retrieval to match $top_{k}$ similar edges $\langle v_{i},e_{i,j},v_{j}\rangle$, following Equation 1. With each vertex $v_{j}$ from these edges we initialize a context queue $C_{queue}$ for breadth-first local search (Voudouris et al., 2010).
@@
To fully exploit the logical relations over the graph and hop from indirectly relevant vertices to relevant ones, we introduce breadth-first local search which utilizes the LLM to choose the most appropriate neighbor for each $v_{j}$ in $C_{queue}$ to append to the tail of the queue. Specifically, for each $v_{j}$ in $C_{queue}$ in each round of hop, we leverage LLM to reason over all the questions from its out edges to choose one $e_{j,k}$ with the question which the LLM regards as the most helpful for answering $q$ and append vertex $v_{k}$ to $C_{queue}$. After hopping from each vertex in the current $C_{queue}$ we can expand the context with at most $top_{k}$ new vertices. From these new vertices we continue the next round of hop. Since different vertices may hop to the same vertex, we believe the vertices with more visits are more important for answering $q$, and use a counter $C_{count}$ to track the number of visits for each vertex and measure its importance. By conducting $n_{hop}$ rounds of hop, we realize reasoning-augmented graph traversal and expand the context length to at most $(n_{hop}+1)\times top_{k}$.
@@
To avoid including too many intermediate vertices during the traversal, we introduce a novel metric Helpfulness $H(\cdot)$ that integrates similarity and logic to re-rank and then prune the traversal counter $C_{count}$. We calculate $H_{i}$ following Equation 2 for each $v_{i}$ in $C_{count}$ and keep the $top_{k}$ vertices with the highest $H_{i}$, where hybrid textual similarity $\text{SIM}(v_{i},q)$ calculates the average lexical and semantic similarity between the passage in $v_{i}$ and query $q$ following Equation 1; and $\text{IMP}(v_{i},C_{count})$ is defined as the normalized number of visits of $v_{i}$ in $C_{count}$ during traversal following Equation 3. We prune $C_{count}$ by retaining $top_{k}$ vertices with the highest $H$ value, resulting in the final context $C$.
$H_{i}=\frac{\text{SIM}(v_{i},q)+\text{IMP}(v_{i},C_{count})}{2}$ (2)
$\text{IMP}(v_{i},C_{count})=\frac{C_{count}[v_{i}]}{\sum_{v_{j}\in C_{count}}C_{count}[v_{j}]}$ (3)
Input: $q$, $top_{k}$, $n_{hop}$, $G$
@@ 1 $\mathbf{v}_{q}$leftarrow EMB($q$);
3 $C_{queue}\leftarrow$ Retrieve($\mathbf{v}_{q}$, $k_{q}$, $G$);
4 $C_{count}\leftarrow$Counter$(C_{queue})$;
@@ 4.1 Experimental Setups
We collect several multi-hop QA datasets to evaluate the performance of HopRAG. We use HotpotQA dataset (Yang et al., 2018), 2WikiMultiHopQA dataset (Ho et al., 2020) and MuSiQue dataset (Trivedi et al., 2022). Following the same procedure as (Zhang et al., 2024), we obtain 1000 questions from each validation set of these three datasets. See Appendix A.2 for details.
@@
We compare HopRAG with a variety of baselines: (1) unstructured RAG - sparse retriever BM25 (Robertson and Zaragoza, 2009a) (2) unstructured RAG - dense retriever BGE (Xiao et al., 2024; Karpukhin et al., 2020) (3) unstructured RAG - dense retriever BGE with query decomposition (Min et al., 2019) (4) unstructured RAG - dense retriever BGE with reranking (Nogueira and Cho, 2020) (5) tree-structured RAG - RAPTOR (Sarthi et al., 2024) (6) tree-structured RAG - SiReRAG (Zhang et al., 2024) (7) graph-structured RAG - GraphRAG (Edge et al., 2024a) with the local search function (8) graph-structured RAG - HippoRAG (Gutiérrez et al., 2025). For structured RAG baselines, we follow the same setting as previous work (Zhang et al., 2024).
@@
To measure the answer quality of different methods, we adopt exact match (EM) and F1 score which focus on the accuracy between a generated answer and the corresponding ground truth. We also use retrieval metrics to compare graph-based methods. Since tree-based methods like SiReRAG (Zhang et al., 2024) and RAPTOR (Sarthi et al., 2024) create new candidates (e.g., summary nodes) in the retrieval pool, it would be unfair to use retrieval metrics to compare them with others. We report both the answer and retrieval metrics in the ablations and discussion on HopRAG. See Appendix A.3 for more metric details.
@@
We use BGE embedding model for semantic vectors at 768 dimensions. To avoid the loss of semantic information caused by chunking at a fixed size, we adopt the same chunking methods utilized in the original datasets respectively. GPT-4o-mini serves as both the model generating in-coming and out-coming questions when constructing the graph index, and the reasoning model for graph traversal. We use two reader models GPT-4o and GPT-3.5-turbo to generate the response given the context with 20 retrieval candidates and $n_{hop}=4$. See Appendix A.4 for more setting details.